OpenStudy (anonymous):

prove that for every real number e > 0, there is a natural number M such that if m>n>M, then 1/n - 1/m < e

OpenStudy (anonymous):

Looks like you're trying to show that the sequence \(a_n=\dfrac{1}{n}\) is a Cauchy sequence?

OpenStudy (anonymous):

im honestly not sure, but as a hint in the textbook is says "because m is positive, the statement 1/n - 1/m < e follows from 1/n < e" and when by e it says epislion in the book

OpenStudy (anonymous):

Letting \(M=\dfrac{1}{\epsilon}\) should work.

OpenStudy (anonymous):

In doing so, you have \[\begin{align*} \left|\frac{1}{n}-\frac{1}{m}\right|&=\left|\frac{m-n}{mn}\right|\\ &=\left|\frac{n-m}{mn}\right|&\text{by a property of absolute value}\\ &=\frac{n-m}{mn}&\text{because }m,n>0\\ &<\frac{n}{mn}\\ &=\frac{1}{m}\\ &<\frac{1}{N}&\text{since }m>N\\\\ &=\epsilon \end{align*}\]

OpenStudy (anonymous):

Sorry, that should be an \(M\), not an \(N\)

OpenStudy (anonymous):

Okay Thanks, im trying to process this lol.

OpenStudy (anonymous):

Is it that some \(n,m\) exist? Or is there something else about \(n\) and \(m\) we should know?

OpenStudy (anonymous):

Usually, if you're working this kind of proof out on paper, you'd start by trying to find a proper \(M\) that would work. Think of it in the way you'd write an \(\epsilon-\delta\) proof for a limit.

OpenStudy (anonymous):

what i typed in the original response is all that is said in the textbook question and that hint was in the back of the book

OpenStudy (anonymous):

So \(n\) and \(m\) don't have to be integers?

OpenStudy (anonymous):

so just basically what u did was combined 1/n - 1/m and then since we know that m > n > M, it shows that when combined it is less than the n/nm ( aka 1/m ) and 1/m is less than 1/M

OpenStudy (anonymous):

@wio i guess

OpenStudy (anonymous):

The fact that \(\epsilon\) is a real number leads me to believe \(m\) and \(n\) are at least positive real numbers. Obviously, they can't be zero, and a proof for negative \(m,n\) would be similar, I think. And yes @Sobie1337, that's the gist of it.

OpenStudy (anonymous):

okay thanks so much. Proofs confuse me so much.... i dont even know why im taking a proof class.

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

shouldn't it be (m-n)/(mn) < m/(mn) = 1/n? because (n-m)/(mn) < 0